### v2

# Khovanov homology can distinguish exotic Mazur manifolds

Abstract.
A Mazur manifold is a compact, contractible $4$-manifold that has a handle decomposition with a single $1$-handle and a single $2$-handle. We show that Khovanov homology can distinguish certain exotic Mazur manifolds.
The author was partially supported by the ILJU Academy and Culture Foundation, the Simons collaboration New structures in low-dimensional topology, and a Princeton Centennial Fellowship.
1. Introduction
Ren and Willis [RW24] gave the first analysis-free proof of the existence of exotic compact, orientable $4$-manifolds; their main tools are the Khovanov skein lasagna module and the Khovanov skein lasagna $s$-invariant. We first deduce from Hayden and Sundberg’s theorems [HS24] that Khovanov homology can distinguish exotic compact, orientable $4$-manifolds (Corollary 2.3). Then, we use similar ideas to show that Khovanov homology can distinguish exotic Mazur manifolds, i.e. compact, contractible $4$-manifolds that have handle decompositions with a single $1$-handle and a single $2$-handle. The first pair of exotic compact, contractible $4$-manifolds was obtained by Akbulut and Ruberman [AR16], and examples of exotic pairs of Mazur manifolds have previously been discovered as well [AY19, HMP21].
Theorem 1.1.
For every integer $k\geq 1$, Khovanov homology can distinguish the exotic pair of Mazur manifolds of Figure 1.1.
Figure 1.1. Exotic pairs of Mazur manifolds ($k\in\mathbb{Z}$, $k\geq 1$).
Theorem 1.1 is proved in Section 4. These $4$-manifolds are the exteriors of the disks in $\overline{\mathbb{CP}^{2}}\backslash{\rm int}D^{4}$ given by blowing up Hayden and Sundberg’s exotic asymmetric disks in $D^{4}$ (Figure 2.1) at a point on the disks: see Figure 4.1 for alternative handle diagrams of these manifolds. We thank Kyle Hayden for pointing out that they are Mazur manifolds.
Ren and Willis [RW24, Section 6.11] defined the Khovanov cobordism map for oriented surfaces in $\mathbb{CP}^{2}\backslash{\rm int}D^{4}$, and showed, using skein lasagna modules, that it only depends on the isotopy class of the surface rel $\partial$. A key ingredient of our proof of Theorem 1.1 is Lemma 3.3, which says that this cobordism map is invariant under diffeomorphisms of $\mathbb{CP}^{2}\backslash{\rm int}D^{4}$ that fix the boundary pointwise. Although we were motivated by skein lasagna modules, we also present a proof in Subsection 3.1 that avoids the theory of skein lasagna modules, with the aim of making the argument more transparent.
Theorem 1.1 still leaves the following questions open.
Question 1.2.
Can Khovanov homology distinguish exotic closed, oriented $4$-manifolds?
Question 1.3.
Can Khovanov homology distinguish exotic closed, oriented, simply connected $4$-manifolds?
Conventions
Homeomorphisms, diffeomorphisms, and isotopies are rel $\partial$ if they fix the boundary pointwise.
If $M$ is an oriented manifold, then M-M is $M$ equipped with the opposite orientation.
The oriented boundary of the standard $4$-ball $D^{4}$ is $S^{3}$, and if $\Sigma\subset D^{4}$ is a properly embedded surface with boundary $L\subset S^{3}$, then the corresponding cobordism maps on Khovanov homology are
$Kh(\Sigma):Kh(\emptyset)\to Kh(L),\ Kh(m(L))\to Kh(\emptyset),$
where $m(L)$ is the mirror of $L$. Similarly, if we denote the mirror of $\Sigma$ as $m(\Sigma)$, then it induces
$Kh(m(\Sigma)):Kh(\emptyset)\to Kh(m(L)),\ Kh(L)\to Kh(\emptyset).$
Remark 1.4.
Our orientation conventions, in effect, are the same as [MWW22] (see [MWW22, Example 5.6]) and [RW24, Section 6.11], but are the opposite of [HS24].
Acknowledgements
We thank Peter Ozsváth for his continuous support and helpful discussions. The main observation of this work was made while rereading the author’s Part III essay during a visit to Duke University. We thank Duke University for their hospitality and Jacob Rasmussen for advising the author’s Part III essay Khovanov Homology and Embedded Surfaces. We thank Nathan Dunfield for help with SnapPy, Kyle Hayden, Adam Levine, Lisa Piccirillo, Misha Schmalian, Alison Tatsuoka, and Qiuyu Ren for helpful discussions, and Kyle Hayden and Robert Lipshitz for helpful comments on earlier drafts.
2. Exotic compact, orientable $4$-manifolds from exotic disks
Figure 2.1. [HS24, Figure 6] The exotic slice disks $\Sigma_{k},\Sigma_{k}^{\prime}$ of $J_{k}$
In [HS24, Example 3.6], Hayden and Sundberg define knots $J_{k}\subset S^{3}$ for $k\in\mathbb{Z}$ and a pair of ribbon disks $\Sigma_{k},\Sigma_{k}^{\prime}\subset D^{4}$ that $J_{k}$ bounds (denoted as $J_{m},\Sigma_{m},\Sigma_{m}^{\prime}$ respectively in [HS24]). With respect to the radial height function of $D^{4}$, these ribbon disks have two index $1$ critical points and three index 0 critical points. The knot $J_{k}$ and the two bands that correspond to the index $1$ critical points are drawn in Figure 2.1.
Theorem 2.1 ([HS24, Example 3.6] and proof of [HS24, Theorem 1.1]).
For every integer $k\geq 1$, the ribbon disks $\Sigma_{k},\Sigma_{k}^{\prime}\subset D^{4}$ of Figure 2.1 are topologically ambiently isotopic rel $\partial$ but not smoothly ambiently isotopic rel $\partial$. In fact, there exists an element $\phi\in Kh(J_{k})$ such that $Kh(m(\Sigma_{k}))(\phi)=\pm 1$ and $Kh(m(\Sigma_{k}^{\prime}))(\phi)=0$, where
$Kh(m(\Sigma_{k})),Kh(m(\Sigma_{k}^{\prime})):Kh(J_{k})\to\mathbb{Z}$
are the cobordism maps on Khovanov homology over $\mathbb{Z}$, where $m(\Sigma_{k})$ (resp. $m(\Sigma_{k}^{\prime})$) means the mirror of $\Sigma_{k}$ (resp. $\Sigma_{k}^{\prime}$).
Recall that the cobordism maps on Khovanov homology are well-defined up to sign, and only depend on the isotopy class rel $\partial$ of the surface [Kho00, Jac04, MWW22].
Figure $2.$ Handle diagrams of the exotic pairs of disk exteriors $D^{4}\backslash N(\Sigma_{k})$ and $D^{4}\backslash N(\Sigma_{k}^{\prime})$
We deduce from Theorem 2.1 (Corollary 2.3) that for every $k\geq 1$, the compact, orientable $4$-manifolds of Figure 2.2 are exotic. This follows from (1) that they are the disk exteriors $D^{4}\backslash N(\Sigma_{k})$ and $D^{4}\backslash N(\Sigma_{k}^{\prime})$ (see [GS99, Section 6.2]) where $N$ means open tubular neighborhood, (2) the following well known lemma (Lemma 2.2), and (3) that the mapping class group of the boundary $S_{0}^{3}(J_{k})$ (the 0-surgery of $S^{3}$ along the knot $J_{k}$) of $D^{4}\backslash N(\Sigma_{k})$ and $D^{4}\backslash N(\Sigma_{k}^{\prime})$ are trivial.
Lemma 2.2 ([MWW22, Lemma 4.7], [LS22, Lemma 4.7], [ha]).
Let $\Sigma,\Sigma^{\prime}\subset D^{4}$ be properly embedded surfaces with the same boundary. If the pairs $(D^{4},\Sigma)$ and $(D^{4},\Sigma^{\prime})$ are diffeomorphic rel $\partial$, then $\Sigma$ and $\Sigma^{\prime}$ are smoothly ambiently isotopic rel $\partial$.
Proof.
Assume that there exists a diffeomorphism $f:(D^{4},\Sigma)\to(D^{4},\Sigma^{\prime})$ that is the identity on $\partial D^{4}$. After modifying $f,\Sigma,\Sigma^{\prime}$ if necessary, we can further assume that $\Sigma,\Sigma^{\prime}$ avoid a neighborhood of the origin, say $\frac{1}{4}D^{4}$, and that $f$ is the identity on a neighborhood of the boundary, say $D^{4}\backslash\frac{1}{3}D^{4}$.
Consider a radial smooth ambient isotopy rel $\partial$ $(\varphi_{t}):D^{4}\times I\to D^{4}$ that “dilates” $\frac{1}{4}D^{4}$ to $\frac{1}{2}D^{4}$. Then $(\varphi_{t})$ is a smooth ambient isotopy between $\Sigma$ and $\varphi_{1}(\Sigma)$, and $(f\circ\varphi_{t}\circ f^{-1})$ is a smooth ambient isotopy between $\Sigma^{\prime}$ and $f\circ\varphi_{1}\circ f^{-1}(\Sigma^{\prime})=\varphi_{1}(\Sigma)$. Hence $\Sigma$ and $\Sigma^{\prime}$ are smoothly ambiently isotopic. ∎
Corollary 2.3.
For every integer $k\geq 1$, the compact, orientable $4$-manifolds $D^{4}\backslash N(\Sigma_{k})$ and $D^{4}\backslash N(\Sigma_{k}^{\prime})$ are homeomorphic but not diffeomorphic.
Proof.
First, $D^{4}\backslash N(\Sigma_{k})$ and $D^{4}\backslash N(\Sigma_{k}^{\prime})$ are homeomorphic rel $\partial$ since the disks $\Sigma_{k}$ and $\Sigma_{k}^{\prime}$ are topologically ambiently isotopic rel $\partial$ by Theorem 2.1 (see, for instance, [CP21, Lemma 2.5]). [^1] In fact, that the disks are topologically ambiently isotopic rel $\partial$ is proved by first proving that their exteriors are homeomorphic rel $\partial$: Hayden and Sundberg [HS24] use [CP21] to show that the disks are topologically ambiently isotopic, and Conway and Powell [CP21] first show that the disk exteriors are homeomorphic rel $\partial$ (see the last paragraph of the proof of Theorem 1.4).
Let us show that $D^{4}\backslash N(\Sigma_{k})$ and $D^{4}\backslash N(\Sigma_{k}^{\prime})$ are not diffeomorphic. First, Theorem 2.1 and Lemma 2.2 imply that $(D^{4},\Sigma_{k})$ and $(D^{4},\Sigma_{k}^{\prime})$ are not diffeomorphic rel $\partial D^{4}$, and so $D^{4}\backslash N(\Sigma_{k})$ and $D^{4}\backslash N(\Sigma_{k}^{\prime})$ are not diffeomorphic rel $\partial$. Hence, we are left to show that the mapping class group of the boundary $S_{0}^{3}(J_{k})$ is trivial for all $k\geq 1$. (To the cautious reader: the below indeed also checks that there are no orientation reversing self-diffeomorphisms of $S_{0}^{3}(J_{k})$.)
To show that ${\rm MCG}(S_{0}^{3}(J_{k}))$ is trivial, we use SnapPy [CDGW] inside Sage [The25]. Let us first check it for $k=1$; we handle the general case in Appendix A similarly to [HS24, Proposition A.3 (a)]. First, input $J_{1}$ [Nah25, j1.lnk] as a snappy.ManifoldHP object called L. The following code verifies that $S_{0}^{3}(J_{1})$ is hyperbolic and finds all geodesics with length $\leq 1$.
L.dehn_fill((0,1))
L.verify_hyperbolicity() # True
L.length_spectrum_alt(max_len=1, verified=True, bits_prec=1000)
SnapPy outputs a unique geodesic with real length $0.92213444882961\cdots$ and word cJQpeID (SnapPy may output a different word). Hence, any isometry of $S_{0}^{3}(J_{1})$ must fix this geodesic. Now, use the following code to check that the isometry group of the cusped hyperbolic manifold given by drilling out that geodesic is trivial.
R = L.drill_word(’cJQpeID’,verified=True).filled_triangulation().canonical_retriangulation(verified=True)
len(R.isomorphisms_to(R)) # 1
3. Khovanov homology and oriented surfaces in $\mathbb{CP}^{2}\backslash{\rm int}D^{4}$
For properly embedded, oriented surfaces $S$ in $k\mathbb{CP}^{2}\backslash({\rm int}D^{4}\sqcup{\rm int}D^{4})$, Ren and Willis define [RW24, Section 6.11] a map on Khovanov homology, which is well-defined up to sign and only depends on the isotopy class rel $\partial$ of $S$. In this section, we recall the definition for the special case where $S$ is in $D^{4}\#\mathbb{CP}^{2}=:(\mathbb{CP}^{2})^{\circ}$, [^2] We write $D^{4}\#\mathbb{CP}^{2}$ to make explicit that its oriented boundary is $S^{3}$ and to avoid confusion in Section 4, where we blow up surfaces in $D^{4}$. sketch a direct (skein lasagna free) proof that it only depends on the isotopy class rel $\partial$, and show that the map is invariant under diffeomorphisms of $(\mathbb{CP}^{2})^{\circ}$ rel $\partial$. The readers who are (resp. are not) familiar with skein lasagna modules may skip Subsection 3.1 (resp. Subsection 3.2).
3.1. A direct argument
Let us first recall the definition of the cobordism map on Khovanov homology for oriented surfaces $S$ properly embedded in $(\mathbb{CP}^{2})^{\circ}$. For simplicity, we work over $\mathbb{Z}$. Further assume that $S$ intersects the core $\mathbb{CP}^{1}$ transversely, positively at $p$ points and negatively at $q$ points. Let $N(\mathbb{CP}^{1})$ be a small open tubular neighborhood of $\mathbb{CP}^{1}$. Then, $S\cap\partial N(\mathbb{CP}^{1})\subset\partial N(\mathbb{CP}^{1})\cong S^{3}$ is the negative $(p+q,p+q)$ torus link, where $p$ of the strands are oriented oppositely to the other $q$ strands. We denote the mirror of such oriented links as $T(p+q,p+q)_{p,q}$. The cobordism map $Kh_{\mathbb{CP}^{2}}(S)$ [^3] We denote it as $Kh_{\mathbb{CP}^{2}}(S)$ for notational clarity and to emphasize that $S$ is in $(\mathbb{CP}^{2})^{\circ}$. is defined, up to sign, as the composition
(3.1) $Kh_{\mathbb{CP}^{2}}(S):Kh(m(\partial S))\to Kh(T(p+q,p+q)_{p,q})\to\mathbb{Z},$
where we define the two maps as follows.
[…]
Let ${\rm gr}_{q}(p,q)$ be the quantum filtration degree of the Lee generator of $T(p+q,p+q)_{p,q}$ over $\mathbb{Q}$. Then, Ren shows [Ren24, Corollary 2.2] that
$Kh^{0,{\rm gr}_{q}(p,q)}(T(p+q,p+q)_{p,q})\cong\mathbb{Z}.$
The second map of (3.1) is projection onto the $(0,{\rm gr}_{q}(p,q))$ grading summand.
Remark 3.1.
Ren [Ren24, Theorem 1.1] also shows that ${\rm gr}_{q}(p,q)=(p-q)^{2}-2{\rm max}(p,q)$, and so the cobordism map $Kh_{\mathbb{CP}^{2}}(S)$ has $(h,q)$-bidegree $(0,\chi(S)-\alpha^{2}+|\alpha|)$, where $\mathbb{Z}$ is supported in bidegree $(0,0)$ and $S]\in H_{2}((\mathbb{CP}^{2})^{\circ})$ is $\alpha=p-q$ times a generator.
Remark 3.2.
From this description, we can see that if $S$ intersects the core $\mathbb{CP}^{1}$ of $\mathbb{CP}^{2}$ exactly once, then the cobordism map $Kh_{\mathbb{CP}^{2}}(S)$ is the cobordism map of the surface in $D^{4}$ given by blowing down $\mathbb{CP}^{1}$. In other words: let $\Sigma\subset D^{4}$ be the surface obtained from $S$ by replacing $\overline{N}(\mathbb{CP}^{1})$ with $D^{4}$ and capping the unknot $S\cap\partial N(\mathbb{CP}^{1})$ with a boundary parallel disk in this $D^{4}$. Then $Kh_{\mathbb{CP}^{2}}(S)$ is the cobordism map $Kh(\Sigma):Kh(m(\partial S))\to\mathbb{Z}$ induced by $\Sigma$.
Let us sketch a direct proof that $Kh_{\mathbb{CP}^{2}}(S)$ only depends on the isotopy class rel $\partial$ of $S$ up to sign. This is similar to the proof that $\Phi^{-1}$ is well-defined in the proof of [MN22, Theorem 1.1]. Since the Khovanov cobordism map only depends, up to sign, on the isotopy class rel $\partial$ for surfaces in $0,1]\times S^{3}$, we only have to check that $Kh_{\mathbb{CP}^{2}}(S)$ is invariant, up to sign, under isotopies supported near $\mathbb{CP}^{1}$, and hence only for the following two kinds of isotopies: (1) those that are induced by moving the intersections with $\mathbb{CP}^{1}$ around, and (2) those that introduce or remove two intersections with $\mathbb{CP}^{1}$, of opposite sign.
The first case corresponds to checking that isotoping the strands of $T(p+q,p+q)_{p,q}$ induces an isomorphism on $Kh^{0,{\rm gr}_{q}(p,q)}\cong\mathbb{Z}$, which is clear.
For the second case, consider the composition of two saddle cobordisms
(3.2) $T(p+q,p+q)_{p,q}\to T(p+q,p+q)_{p,q}\sqcup U\to T(p+q+2,p+q+2)_{p+1,q+1}$
where $U$ is an unlinked unknot. The second case corresponds to checking that the induced map
(3.3) $Kh^{0,{\rm gr}_{q}(p,q)}(T(p+q,p+q)_{p,q})\to Kh^{0,{\rm gr}_{q}(p+1,q+1)}(T(p+q+2,p+q+2)_{p+1,q+1})$
is an isomorphism. This follows from [Ren24, Theorem 2.1 (1)]: the first statement of [Ren24, Theorem 2.1 (1)] is that $Kh^{0,i}(T(p+q,p+q)_{p,q})=0$ for $i<{\rm gr}_{q}(p,q)$. Hence,
$Kh^{0,{\rm gr}_{q}(p,q)}(T(p+q,p+q)_{p,q})$ $=Kh^{0,{\rm gr}_{q}(p,q)}(T(p+q,p+q)_{p,q})\oplus Kh^{0,{\rm gr}_{q}(p,q)-2}(T(p+q,p+q)_{p,q})$
$\cong Kh^{0,{\rm gr}_{q}(p,q)-1}(T(p+q,p+q)_{p,q}\sqcup U),$
and the first saddle cobordism of (3.2) induces this isomorphism. The second statement of [Ren24, Theorem 2.1 (1)] is that the second saddle cobordism of (3.2) induces an isomorphism
$Kh^{0,{\rm gr}_{q}(p,q)-1}(T(p+q,p+q)_{p,q}\sqcup U)\to Kh^{0,{\rm gr}_{q}(p+1,q+1)}(T(p+q+2,p+q+2)_{p+1,q+1}).$
Combining the above two, we obtain that (3.3) is an isomorphism.
Now, we show that $Kh_{\mathbb{CP}^{2}}(S)$ is invariant under diffeomorphisms of $(\mathbb{CP}^{2})^{\circ}$ rel $\partial$.
Lemma 3.3.
Let $S,S^{\prime}$ be properly embedded, oriented surfaces in $(\mathbb{CP}^{2})^{\circ}:=D^{4}\#\mathbb{CP}^{2}$ such that $S]=[S^{\prime}]\in H_{2}((\mathbb{CP}^{2})^{\circ};\mathbb{Z})$. If there exists a diffeomorphism
$((\mathbb{CP}^{2})^{\circ},S)\cong((\mathbb{CP}^{2})^{\circ},S^{\prime})\ {\rm rel}\ \partial,$
then the induced maps
$Kh_{\mathbb{CP}^{2}}(S),Kh_{\mathbb{CP}^{2}}(S^{\prime}):Kh(m(\partial S))\to\mathbb{Z}$
agree up to sign.
Proof.
We proceed similarly to the proof of Lemma 2.2. View $(\mathbb{CP}^{2})^{\circ}$ as a unit disk bundle over the core $\mathbb{CP}^{1}$. For $r\in(0,1]$, denote $N_{r}\subset(\mathbb{CP}^{2})^{\circ}$ as the open disk subbundle over $\mathbb{CP}^{1}$ with radius $r$.
Assume that there exists a diffeomorphism rel $\partial$ $f:((\mathbb{CP}^{2})^{\circ},S)\to((\mathbb{CP}^{2})^{\circ},S^{\prime})$. After isotoping $f,S,S^{\prime}$ if necessary, assume that $S\cap N_{1/4}$ and $S^{\prime}\cap N_{1/4}$ are a finite number of fiber disks, and that $f$ is the identity on $(\mathbb{CP}^{2})^{\circ}\backslash N_{1/3}$. Consider a fiberwise smooth ambient isotopy rel $\partial$ $(\varphi_{t}):(\mathbb{CP}^{2})^{\circ}\times I\to(\mathbb{CP}^{2})^{\circ}$ that “dilates” $N_{1/4}$ to $N_{1/2}$. Then $(\varphi_{t})$ is a smooth ambient isotopy between $S$ and $\varphi_{1}(S)$, and $(f\circ\varphi_{t}\circ f^{-1})$ is a smooth ambient isotopy between $S^{\prime}$ and $f\circ\varphi_{1}(S)$.
We claim that the cobordism maps $Kh_{\mathbb{CP}^{2}}(\varphi_{1}(S))$ and $Kh_{\mathbb{CP}^{2}}(f\circ\varphi_{1}(S))$ agree. Note that the surfaces $\varphi_{1}(S)$ and $f\circ\varphi_{1}(S)$ agree on $(\mathbb{CP}^{2})^{\circ}\backslash N_{1/3}$. Let $\varphi_{1}(S)$ intersect $\mathbb{CP}^{1}$ at $p$ points positively and $q$ points negatively. Since $\varphi_{1}(S)\cap N_{1/3}$ is some number of fiber disks, $Kh_{\mathbb{CP}^{2}}(\varphi_{1}(S))$ is induced by the Khovanov map given by the cobordism
$((\mathbb{CP}^{2})^{\circ}\backslash N_{1/3},\varphi_{1}(S)\cap((\mathbb{CP}^{2})^{\circ}\backslash N_{1/3})):(S^{3},m(\partial S))\to(S^{3},T(p+q,p+q)_{p,q}).$
By modifying $f\circ\varphi_{1}(S)$ near $\mathbb{CP}^{1}$ if necessary, assume that there exists an $\varepsilon>0$ such that $(f\circ\varphi_{1}(S))\cap N_{\varepsilon}$ is $(p+q+2\ell)$ many fiber disks. Note that $\ell\geq 0$. Then, we are left to show that
(3.4) $(\overline{N_{1/3}}\backslash N_{\varepsilon},S\cap(\overline{N_{1/3}}\backslash N_{\varepsilon})):(S^{3},T(p+q,p+q)_{p,q})\to(S^{3},T(p+q+2\ell,p+q+2\ell)_{p+\ell,q+\ell})$
induces an isomorphism
(3.5) $Kh^{0,{\rm gr}_{q}(p,q)}(T(p+q,p+q)_{p,q})\to Kh^{0,{\rm gr}_{q}(p+\ell,q+\ell)}(T(p+q+2\ell,p+q+2\ell)_{p+\ell,q+\ell}).$
In fact, we claim that if $A$ is a surface in $0,1]\times S^{3}$ with the same domain and codomain as (3.4) and is topologically the disjoint union of $(p+q)$ many annuli with $2\ell$ disks removed, then the induced cobordism map (3.5) is an isomorphism. Note that the surface in (3.4) satisfies this condition. To show the claim, we use cobordism maps on the Lee spectral sequence [Lee05, Ras05, Ras10] over $\mathbb{Q}$. Let $\mathfrak{s}_{p,q}$ (resp. $\mathfrak{s}_{p+\ell,q+\ell}$) be the Lee generator of $T(p+q,p+q)_{p,q}$ (resp. $T(p+q+2\ell,p+q+2\ell)_{p+\ell,q+\ell}$), and let us also denote $\mathfrak{s}_{p,q}$ (resp. $\mathfrak{s}_{p+\ell,q+\ell}$) as its image in the $E_{\infty}$ page of the Lee spectral sequence. Then, for any such $A$, the induced cobordism map on the $E_{\infty}$ page of the Lee spectral sequence over $\mathbb{Q}$ maps $\mathfrak{s}_{p,q}$ to $\pm C\mathfrak{s}_{p+\ell,q+\ell}$, where $C$ only depends on $p,q,\ell$. Since the domain and the codomain of (3.3) are $\mathbb{Z}$ and these gradings are where the Lee generators $\mathfrak{s}_{p,q}$ and $\mathfrak{s}_{p+\ell,q+\ell}$ (respectively) live over $\mathbb{Q}$, the domain and codomain of (3.3) survive to the $E_{\infty}$ page of the Lee spectral sequence over $\mathbb{Q}$. Hence, any such $A$ and $A^{\prime}$ induce the same map (3.5) up to sign. To show that (3.5) is an isomorphism over $\mathbb{Z}$, we only have to check it for one such $A$: one choice is the composition of $\ell$ many of (3.3). ∎
3.2. A skein lasagna argument
In this subsection, we present a skein lasagna proof of Lemma 3.3. For simplicity, we work over $\mathbb{Q}$ in this subsection, unless specified otherwise.
Ren and Willis [RW24, Section 6.11] give a description of the cobordism map
$Kh_{\mathbb{CP}^{2}}(S):Kh(m(\partial S))\to\mathbb{Q}$
for oriented surfaces $S\subset(\mathbb{CP}^{2})^{\circ}$, in terms of the Khovanov-Rozansky $\mathfrak{gl}_{2}$ skein lasagna module. Note that this latter description is phrased in terms of the $\mathfrak{gl}_{2}$ homology $KhR_{2}$ instead of $Kh$, which can be thought as a renormalization of $Kh$ of the mirror of the link [RW24, Equation (8)]; hence the cobordism map is defined for oriented surfaces in $D^{4}\#\overline{\mathbb{CP}^{2}}$. For a quick introduction to the $\mathfrak{gl}_{2}$ skein lasagna modules, we refer the readers to [RW24, Sections 2.1-2.4]; see the original paper [MWW22] for the general setting.
Let $S$ be an oriented surface in $D^{4}\#\overline{\mathbb{CP}^{2}}$. Let us recall the skein lasagna description of the cobordism map
(3.6) $KhR_{2,\overline{\mathbb{CP}^{2}}}(S):KhR_{2}(m(\partial S))\to\mathbb{Q}.$
Let $\alpha:=[S]\in H_{2}(D^{4}\#\overline{\mathbb{CP}^{2}})=H_{2}(\overline{\mathbb{CP}^{2}}).$ Ren and Willis show that ${\mathcal{}S}_{0,0,|\alpha|}^{2}(\overline{\mathbb{CP}^{2}};\alpha;\mathbb{Z})\cong\mathbb{Z}$, and define the canonical dual lasagna generator $\theta_{\alpha}\in({\mathcal{}S}_{0}^{2}(\overline{\mathbb{CP}^{2}};\alpha;\mathbb{Q}))_{0,-|\alpha|}^{\ast}$ up to sign, as the dual element that maps a generator of ${\mathcal{}S}_{0,0,|\alpha|}^{2}(\overline{\mathbb{CP}^{2}};\alpha;\mathbb{Z})\cong\mathbb{Z}$ to $1$. Consider the lasagna filling with no input balls given by the surface $S$. This defines an element in ${\mathcal{}S}_{0}^{2}(D^{4}\#\overline{\mathbb{CP}^{2}};\partial S;\alpha)$; denote it also as $S$ by abuse of notation. Then, the cobordism map (3.6) is the image of $S\in{\mathcal{}S}_{0}^{2}(D^{4}\#\overline{\mathbb{CP}^{2}};\partial S;\alpha)$ under the following composition
${\mathcal{}S}_{0}^{2}(D^{4}\#\overline{\mathbb{CP}^{2}};\partial S;\alpha)\xrightarrow[\cong]{\kappa}{\rm Hom}(KhR_{2}(m(\partial S)),\mathbb{Q})\otimes{\mathcal{}S}_{0}^{2}(\overline{\mathbb{CP}^{2}};\alpha)\xrightarrow{{\rm Id}\otimes\theta_{\alpha}}{\rm Hom}(KhR_{2}(m(\partial S)),\mathbb{Q}),$
where the first map $\kappa$ is the Künneth map for the skein lasagna module [MN22, Theorem 1.4], and the second map is given by evaluating with the canonical dual lasagna generator $\theta_{\alpha}\in({\mathcal{}S}_{0}^{2}(\overline{\mathbb{CP}^{2}};\alpha;\mathbb{Q}))_{0,-|\alpha|}^{\ast}$.
A skein lasagna proof of Lemma 3.3 over $\mathbb{Q}$.
We work with the $\mathfrak{gl}_{2}$ homology $KhR_{2}$ and oriented surfaces $S,S^{\prime}\subset D^{4}\#\overline{\mathbb{CP}^{2}}$ such that $\alpha:=[S]=[S^{\prime}]\in H_{2}(D^{4}\#\overline{\mathbb{CP}^{2}})=H_{2}(\overline{\mathbb{CP}^{2}})$. We show that if there exists a diffeomorphism rel $\partial$
$F:(D^{4}\#\overline{\mathbb{CP}^{2}},S)\to(D^{4}\#\overline{\mathbb{CP}^{2}},S^{\prime}),$
then $KhR_{2,\overline{\mathbb{CP}^{2}}}(S^{\prime})=\pm KhR_{2,\overline{\mathbb{CP}^{2}}}(S)$.
Without loss of generality, we may assume that $F$ fixes a collar neighborhood of the boundary; assume that it fixes a neighborhood of the $D^{4}$ summand. Extend $F$ by the identity to a diffeomorphism $G:\overline{\mathbb{CP}^{2}}\to\overline{\mathbb{CP}^{2}}$. Since $F$ is the identity on a neighborhood of the $D^{4}$ summand, one can check by going through the proof of the Künneth formula [MN22, Theorem 1.4] that the following commutes, where $\mathcal{S}_{0}^{2}(F),\mathcal{S}_{0}^{2}(G)$ are the induced isomorphisms on the $\mathfrak{gl}_{2}$ skein lasagna modules.
$\mathcal{S}_{0}^{2}(D^{4}\#\overline{\mathbb{CP}^{2}};\partial S;\alpha)$${\rm Hom}(KhR_{2}(m(\partial S)),\mathbb{Q})\otimes\mathcal{S}_{0}^{2}(\overline{\mathbb{CP}^{2}};\alpha)$$\mathcal{S}_{0}^{2}(D^{4}\#\overline{\mathbb{CP}^{2}};\partial S^{\prime};\alpha)$${\rm Hom}(KhR_{2}(m(\partial S^{\prime})),\mathbb{Q})\otimes\mathcal{S}_{0}^{2}(\overline{\mathbb{CP}^{2}};\alpha)$$\scriptstyle{\kappa}\scriptstyle{\cong}$$\scriptstyle{\mathcal{S}_{0}^{2}(F)}$$\scriptstyle{{\rm{Id}}\otimes\mathcal{S}_{0}^{2}(G)}$$\scriptstyle{\kappa}\scriptstyle{\cong}$
Since ${\mathcal{}S}_{0}^{2}(G)$ is an isomorphism that preserves the bigrading of ${\mathcal{}S}_{0}^{2}(\overline{\mathbb{CP}^{2}};\alpha)$ and ${\mathcal{}S}_{0,0,|\alpha|}^{2}(\overline{\mathbb{CP}^{2}};\alpha;\mathbb{Z})=\mathbb{Z}$, ${\mathcal{}S}_{0}^{2}(G)$ fixes the canonical dual lasagna generator $\theta_{\alpha}\in({\mathcal{}S}_{0}^{2}(\overline{\mathbb{CP}^{2}};\alpha))_{0,-|\alpha|}^{\ast}$ up to sign. Hence,
$KhR_{2,\overline{\mathbb{CP}^{2}}}(F(S))$ $=({\rm Id}\otimes\theta_{\alpha})\circ\kappa(F(S))$
$=({\rm Id}\otimes\theta_{\alpha})\circ\kappa\circ{\mathcal{}S}_{0}^{2}(F)(S)$
$=({\rm Id}\otimes(\theta_{\alpha}\circ{\mathcal{}S}_{0}^{2}(G)))\circ\kappa(S)$
$=\pm({\rm Id}\otimes\theta_{\alpha})\circ\kappa(S)$
$=\pm KhR_{2,\overline{\mathbb{CP}^{2}}}(S).$
4. Exotic Mazur manifolds from Khovanov homology
In this section, we use the Khovanov map for oriented cobordisms in $\mathbb{CP}^{2}$ from Section 3 to show Theorem 1.1. We thank Lisa Piccirillo for suggesting this. Note that we mainly work with the mirrors $m(\Sigma_{k}),m(\Sigma_{k}^{\prime})\subset D^{4}$ because of our orientation conventions (see Remark 1.4 and Theorem 2.1).
Proof of Theorem 1.1.
Let $S_{k}$ (resp. $S_{k}^{\prime}$) be the disk in $D^{4}\#\mathbb{CP}^{2}$ given by blowing up the disk $m(\Sigma_{k})\subset D^{4}$ (resp. $m(\Sigma_{k}^{\prime})$) at a point on the disk. In other words, choose a small closed $4$-ball $B$ such that $(B,B\cap m(\Sigma_{k}))\cong(D^{4},D^{2})$. In particular, $\partial(B\cap m(\Sigma_{k}))\subset\partial B\cong S^{3}$ is an unknot. Replace $(B,B\cap m(\Sigma_{k}))$ by $(\overline{N}(\mathbb{CP}^{1}),F)$, where $\overline{N}(\mathbb{CP}^{1})$ is a closed tubular neighborhood of the core $\mathbb{CP}^{1}$ in $\mathbb{CP}^{2}$, $F$ is a fiber disk, and $\partial F\subset\partial\overline{N}(\mathbb{CP}^{1})\cong S^{3}$ is identified with $\partial(B\cap m(\Sigma_{k}))\subset\partial B$. Let $S_{k}$ be the resulting properly embedded surface in $(\mathbb{CP}^{2})^{\circ}:=D^{4}\#\mathbb{CP}^{2}$. Similarly, let $S_{k}^{\prime}$ be the surface in $(\mathbb{CP}^{2})^{\circ}$ obtained from $m(\Sigma_{k}^{\prime})$ analogously.
Then, by Remark 3.2, the maps
$Kh_{\mathbb{CP}^{2}}(S_{k}),Kh_{\mathbb{CP}^{2}}(S_{k}^{\prime}):Kh(J_{k})\to\mathbb{Z}$
are the same as
$Kh(m(\Sigma_{k})),Kh(m(\Sigma_{k}^{\prime})):Kh(J_{k})\to\mathbb{Z},$
respectively.
Hence, $((\mathbb{CP}^{2})^{\circ},S_{k})$ and $((\mathbb{CP}^{2})^{\circ},S_{k}^{\prime})$ are not diffeomorphic rel $\partial$ by Theorem 2.1 and Lemma 3.3, and so $(\mathbb{CP}^{2})^{\circ}\backslash N(S_{k})$ and $(\mathbb{CP}^{2})^{\circ}\backslash N(S_{k}^{\prime})$ are not diffeomorphic rel $\partial$. Using the same method as the proof of Corollary 2.3, one can show that their boundary has trivial mapping class group for all $k\geq 1$. Note that their boundary is $S_{-1}^{3}(m(J_{k}))\cong-S_{1}^{3}(J_{k})$.
The $4$-manifold $(\mathbb{CP}^{2})^{\circ}\backslash N(S_{k})$ can be obtained from $D^{4}\backslash N(m(\Sigma_{k}))$ by attaching a $1$-framed 22-handle along a meridian of $m(\Sigma_{k})$, and an analogous statement holds for $(\mathbb{CP}^{2})^{\circ}\backslash N(S_{k}^{\prime})$. Hence, the first diagrams of Figure 4.1 are $((\mathbb{CP}^{2})^{\circ}\backslash N(S_{k}))$ and $((\mathbb{CP}^{2})^{\circ}\backslash N(S_{k}^{\prime}))$, and these manifolds are homeomorphic since $D^{4}\backslash N(\Sigma_{k})$ and $D^{4}\backslash N(\Sigma_{k}^{\prime})$ are homeomorphic rel $\partial$ (Corollary 2.3). Finally, Figure 4.1 checks that they can be simplified to the handle diagrams of Figure 1.1. This proves Theorem 1.1. ∎
Figure 4.1. Handle calculus that shows that the two manifolds of Figure 1.1 are diffeomorphic to the disk exteriors $((\mathbb{CP}^{2})^{\circ}\backslash N(S_{k}))$ and $((\mathbb{CP}^{2})^{\circ}\backslash N(S_{k}^{\prime}))$ from the proof of Theorem 1.1. In the final step, the handle diagrams for $k=1$ are shown.
Appendix A Computation of the mapping class group
[…]
L.dehn_fill((0,1),0)
R = L.filled_triangulation().canonical_retriangulation(verified=True)
R.verify_hyperbolicity() # True
len(R.isomorphisms_to(R)) # 1
[…]
Remark A.1.
[…]
Figure A.1. [HS24, Figure 14] A 22-component link $J_{0}\sqcup U$ [Nah25, j0u.lnk].
[…]
R.length_spectrum_alt(max_len=1, verified=True, bits_prec=1000)
Now, R.cusp_areas(verified=True) outputs $13.24\cdots$, and hence we use the following code to find all the slopes on the cusp with length at most $37.5$, which ensures a normalized length of $37.5/\sqrt{13.3}>10.1$.
R.short_slopes(verified=True, length=37.5)
This leaves us to check $k=1,\cdots,17$ in the same way as we checked $k=1$ in Section 2. One can use the following code:
L.dehn_fill((0,1),0)
for k in range(1,18):
L.dehn_fill((1,k),1)
print(L.verify_hyperbolicity()[0])
sp = L.length_spectrum_alt(max_len=1, verified=True, bits_prec=1000)
print(sp)
R = L.drill_word(sp[0].word, verified=True, bits_prec=1000).filled_triangulation().canonical_retriangulation(verified=True)
print("k =",k,"; MCG =",len(R.isomorphisms_to(R)))
Since verified=True, if the above code runs successfully, then SnapPy will have proved, using interval arithmetic, that $S_{0,1/k}^{3}(J_{0},U)$ is hyperbolic and will have provably computed the rank of its mapping class group. However, this does not mean that $S_{0,1/k}^{3}(J_{0},U)$ being hyperbolic and it having a geodesic of length $\leq 1$ always imply that the code will run successfully. Indeed, using the above given link diagram of $J_{0}\sqcup U$ [Nah25, j0u.lnk], SnapPy successfully proved that ${\rm MCG}(S_{0,1/k}^{3}(J_{0},U))$ is trivial for all $k\geq 1$ and that ${\rm MCG}(S_{1,1/k}^{3}(J_{0},U))$ is trivial for all $k\geq 2$, but failed to prove it for $S_{1,1}^{3}(J_{0},U)$. For this case, working directly with the knot diagram of $J_{1}$ [Nah25, j1.lnk] as in Section 2 worked for us.